Account Title Cash Accounts receivable Office supplies Trucks Accumulated depreciation-Trucks Land Accounts payable InterestWIGO VICCL December 31 Assets Liabilities Equity Total equity
Answer #1

A B C 1 2 WILSON TRUCKING COMPANY Income Statement Forf Year Ended December 31 3 $ 122,000 4 Revenue 5 Trucking Fees earned 6$ 6,683 $ 27,783 $ 153,000 $ -31,518 $ 121,482 $ 48,000 $ 169,482 $ 197,265 30 Office Supplies 31 Total Current assets 32 Pla

A B с 1 WILSON TRUCKING COMPANY 2 Income Statement 3 Forf Year Ended December 31 4 Revenue 5 Trucking Fees earned 122000 6 Ex6683 =+C28+C29+C30 153000 |-31518 =+B33+B34 48000 =+C34+C35 =+C31+C36 30 Office Supplies 31 Total Current assets 32 Plant ass